OpenStudy (anonymous):

HELP PLEASE!! ILL GIVE A MEDAL TO WHOEVER ANSWER IT! :) What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

4x-2y=12 -4x 2y=-12 0 = 0

OpenStudy (anonymous):

well 4x-2y = 12, on dividing both sides by 2 is 2x-y = 6. so 2x-y = 6 and 2x-y = 6. both equations are same. hence there is actually no UNIQUE solution to the equation. It can however have infinite solutions if u plug in one value of x then u get the y variable.

OpenStudy (anonymous):

Thank you soo much @luckythebest and @devinthedev